OpenStudy (anonymous):

Find the derivative of V= 1000 ( 1- t/60 ) ^2 using limits

OpenStudy (anonymous):

is this \[V=1000(1-\frac{t}{60})^2\]?

OpenStudy (anonymous):

??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok so you want \[\lim_{h\rightarrow 0}\frac{V(t+h)-V(t)}{h}\] \[1000\lim_{h\rightarrow 0}\frac{1-\frac{t+h}{60})-(1-\frac{t}{60})}{h}\]

OpenStudy (anonymous):

all the work is the algebra in the numerator, so lets do that separately, then divide by h

OpenStudy (anonymous):

yes the 1000 should be behind the limit isnt it?

OpenStudy (anonymous):

\[\lim c f(x)=c\lim f(x)\] so i can pull it right out front

OpenStudy (anonymous):

ohhhh ok

OpenStudy (anonymous):

i did however forget the square, so let me rigth it correctly \[1000\lim_{h\rightarrow 0}\frac{(1-\frac{t+h}{60})^2-(1-\frac{t}{60})^2}{h}\]

OpenStudy (anonymous):

*write

OpenStudy (anonymous):

man this is really going to be a pain with the square, but it is just algebra. again lets focus only on the numerator

OpenStudy (anonymous):

yea, i think i messed the squaring too ..

OpenStudy (anonymous):

messed up*

OpenStudy (anonymous):

\[(1-\frac{t}{60})=1-\frac{t}{30}+\frac{t^2}{3600}\]

OpenStudy (anonymous):

\[(-120+2t)*1000/60^2\]

OpenStudy (anonymous):

where did the t+h go?

OpenStudy (anonymous):

\[(1-\frac{t+h}{60})^2=1-\frac{t}{30}+\frac{t^2}{3600}-\frac{h}{30}+\frac{th}{1800}+\frac{h^2}{3600}\]

OpenStudy (anonymous):

i did the second one first

OpenStudy (anonymous):

because it was easiest

OpenStudy (anonymous):

so now we subtract \[(1+\frac{t+h}{60})^2-(1-\frac{t}{60})^2=-\frac{h}{30}+\frac{th}{1800}+\frac{h^2}{3600}\]

OpenStudy (anonymous):

now we can divide by h

OpenStudy (anonymous):

get \[-\frac{1}{30}+\frac{t}{1800}+\frac{h}{3600}\] let go to zero and get \[\frac{t}{1800}-\frac{1}{30}\] multiply by 1000 and get \[\frac{5t}{9}-\frac{100}{3}\]

OpenStudy (anonymous):

ohh, I caught my mistake now :) thanks

OpenStudy (anonymous):

yw